OpenStudy (anonymous):

If the point P(-3/5,y) lies on the unit circle and P is in the second quadrant, what does y equal?

OpenStudy (anonymous):

positive 4/5 This is because (-3/5)^2 + (4/5)^2 = 1 AND your "y" values have to be positive when your x values are negative to be in the second quadrant.

OpenStudy (anonymous):

Because that is your second quadrant.

OpenStudy (anonymous):

And the equation of your unit circle centered at the origin is: x^2 + y^2 = 1 = 1^2

OpenStudy (anonymous):

So, working with the unknown: (-3/5)^2 + y^2 = 1 and then you solve for y. You get a positive and negative value, but, again, you take the positive value.

OpenStudy (anonymous):

Making sense to you?

OpenStudy (anonymous):

@haleylou56 ?

OpenStudy (anonymous):

yep thanks